Meena earns Tk. 9000 from her poultry firm. Over the day, equipments of egg hatcheries depreciate in value by Tk. 400.Of the remaining amount; Meena gives Tk.250 to the government in sales taxes. Later she retains 290 taka in her business for buying new machines and receives 470 as old scheme benefit. From the remaining amount that Meena takes home, she pays Tk.500 as income taxes. Based on this information, compute Meena’s contribution to the following measures of income:

a. Gross Domestic product

b. Net National Product

c. National Income

d. Personal Income

e. Disposable Personal Income


1
Expert's answer
2021-04-23T07:46:51-0400

a. Gross Domestic product

Tk. 9000

b. Net National Product


"9000-400-250-290-500=7560"

c. National Income


"250+500=750"

d. Personal Income


"9000-400-250-290+470=8530"

e. Disposable Personal Income


"8530-500=8030"